# PeakWeather: MeteoSwiss Weather Station Measurements for Spatiotemporal Deep Learning

### v1

@@ 1 Introduction
Weather forecasts provide essential information for protecting lives and properties, as well as supporting decision-making in everyday activities. The quality of these forecasts has steadily improved over recent decades, driven by advances in numerical weather prediction (NWP) and high-performance computing (HPC) (bauer2015quiet, ). NWP involves using mathematical models that simulate the weather based on current conditions of atmosphere, land surfaces and oceanic conditions. Both deterministic forecasts and ensembles of simulations are produced, with ensemble forecasts offering a probabilistic view of future states and serving as a critical tool for risk-informed decision-making. However, the complexity of the models, combined with the demand for high-resolution predictions in space and time, makes state-of-the-art NWP computationally expensive—often requiring several hours of runtime on modern HPC systems.
@@
In parallel, the rise of deep learning (DL) has brought remarkable progress to a range of fields, including meteorology (lam2023learning, ; bi2022pangu, ; pathak2022fourcastnet, ; price2025probabilistic, ). Beyond breakthroughs in model architectures, this progress has been enabled by the availability of long-term, high-quality reanalysis datasets. Reanalyses combine modern NWP models with historical observational data—from ground stations, satellites, radar, ships, and aircraft—to produce a consistent and comprehensive estimate of past atmospheric states. DL models can learn to emulate the underlying physics captured in these massive datasets, enabling accurate and scalable data-driven weather forecasting. One of the key advantages of such models is their efficiency: once trained, they are significantly cheaper to run operationally than classical physics-based NWP. Nonetheless, these models still rely on NWP-generated analyses at inference time to initialize forecasts.
@@
To overcome the need for NWP-based analysis and reanalysis data altogether, Direct Observation Prediction (DOP) methods have been proposed (mcnally2024data, ). DOP approaches aim to learn directly from raw observational data, eliminating the need for NWP data during both training and inference. When successfully implemented, DOP models can offer faster, more cost-effective forecasting pipelines and avoid potential biases introduced by physics-based simulations. DOP has shown promising results in short-range weather forecasting, particularly for lead times ranging from minutes to a few hours (ravuri2021skilful, ; zhang2023skilful, ), a domain known as nowcasting. A typical nowcasting application is the extrapolation of radar image sequences to predict imminent precipitation. While DOP has proven effective in short time horizons, NWP systems generally maintain an edge in predictive accuracy beyond a few hours.
@@
We introduce PeakWeather, a benchmarking dataset designed to advance deep learning research in spatiotemporal modeling. The dataset consists of high-resolution, ground-based meteorological measurements collected from SwissMetNet (SwissMetNet, )—the network of automatic weather stations operated by the Swiss Federal Office of Meteorology and Climatology, MeteoSwiss. It includes eight key weather variables related to temperature, humidity, precipitation, pressure, sunshine, and wind, recorded every 10 minutes over more than eight years. These observations come from 302 meticulously maintained and quality-validated stations, distributed throughout Switzerland, a country characterized by complex topography, shaped by the Swiss Alps, the Central Plateau, and the Jura Mountains. In addition to meteorological variables, PeakWeather includes topographic descriptors derived from digital elevation models, capturing the unique characteristics of each station’s surrounding area.
@@
To the best of our knowledge, this is the first openly available dataset that combines long-term, high-frequency ground station data with topographic context and state-of-the-art NWP baselines in complex terrain. The complexity of Switzerland’s topography, combined with seasonalities and time variance of the meteorological quantities, pose unique challenges for modeling spatial and temporal dependencies. These factors make PeakWeather an ideal testbed for advancing data-driven methods in a wide array of machine learning tasks, including—but not limited to—time series forecasting, virtual sensing, graph structure learning, and missing data imputation. To support benchmarking, the dataset also includes ensemble forecasts from a state-of-the-art, high-resolution NWP model that has been operational at MeteoSwiss since May 2024. These forecasts serve as a strong physics-based reference point, enabling direct comparisons with machine learning approaches and promoting research in hybrid and alternative forecasting paradigms.
@@
  • Release of PeakWeather dataset (PeakWeatherDataset, ). The dataset contains: (i) A collection of high-quality validated multivariate time series from 302 weather stations comprising eight meteorological variables with a temporal resolution of 10 minutes acquired from January 1st, 2017 to March 31st, 2025; (ii) Topographical descriptors derived from a digital height model describing the Swiss terrain over a 50-meter grid; (iii) Numerical weather predictions from a state-of-the-art high-resolution NWP model.
  • A framework-agnostic, easy-to-use and well-documented Python library (PeakWeatherBaseCode, ) for a seamless interaction with the dataset. The library enables users to easily load, align, and preprocess the data, providing them in ready-to-use tensor formats suitable for deep learning workflows.
  • Evaluation of predictive models for wind forecasting—a task of high practical importance in domains such as renewable energy, aviation, and severe weather management. Indeed, practitioners and researchers can consider and elaborate on different meteorological variables in their research. A range of deep learning architectures, including recurrent neural networks and spatiotemporal graph neural networks, are contrasted on the prediction task against forecasts from a state-of-the-art NWP model. Reported results establish valuable baselines for future research and demonstrate the promising potential of deep learning models as viable alternatives to NWP models in specific weather prediction tasks.
The PeakWeather dataset is released under the CC-BY-4.0 license and the associated library under the BSD-3-Clause. With PeakWeather, we aim at facilitating the use of MeteoSwiss’ data that, starting from Spring 2025, is gradually making available as Open Government Data (OGD) (SwissMetNetDocumentation, ).
@@ 2 Related work
One major category of meteorological datasets consists of gridded reanalysis data produced by NWP models, which assimilate past observations to forecast the weather state. One of the main driving forces behind the recent success of data-driven models is ECMWF Reanalysis v5 (ERA5) (ERA5, ). This dataset contains a detailed estimate of the global state of the atmosphere, land surface and oceans since the 1950s, based on the reanalysis of ECMWF’s Integrated Forecasting System (IFS). Derived from ERA5, WeatherBench (rasp2020weatherbench1, ) and its recent successor WeatherBench2 (rasp2024weatherbench2, ) offer standardized benchmarking datasets and tools, to facilitate the evaluation and comparison of machine learning models for weather forecasting. For the task of statistical postprocessing (vannitsem2021postproc, ), where the goal is to correct systematic errors in NWP forecasts, EUPPBench (demaeyer2023euppbench, ) provides a dataset with aligned IFS forecasts and direct observations for a variety of meteorological variables. So far, however, only coarse-grained NWP data is available in EUPPBench and temporal resolution is limited to 6 hours.
Unlike the datasets above, PeakWeather supports DOP, enabling the development of models that rely solely on past observations and static features, making the final model completely independent from NWP models. While datasets such as WeatherReal (jin2024weatherreal, ), Weather2k (zhu2023weather2k, ), Monash Weather (godahewa2021monash, ), and StationBench (stationbench, ) share this DOP focus, PeakWeather distinguishes itself through its unique combination of characteristics not available in other datasets: quality-controlled data with a high spatial density and 10-minute temporal resolution, spanning over eight years across Switzerland’s challenging topography, and physics-based forecasts from an operational NWP model for comparison.
@@ 3 PeakWeather dataset
We introduce PeakWeather, a publicly available dataset hosted on Hugging Face (PeakWeatherDataset, ). The dataset is organized into three core components: (i) high-resolution historical surface observations, (ii) static topographic features derived from elevation data, and (iii) ensemble forecasts from an operational NWP model. To facilitate easy access and streamlined preprocessing, we also release an open-source Python library (PeakWeatherBaseCode, ) tailored to work seamlessly with the dataset.
@@
Meteorological observations come from the SwissMetNet network, Switzerland’s reference ground-based measurement system for weather and climate monitoring operated by MeteoSwiss. The network in PeakWeather comprises 302 stations in total: 160 standard meteorological stations measuring various weather and climate parameters, complemented by 142 additional stations measuring precipitation. Stations are labeled as either meteorological stations or rain gauges, as shown in Figure 1(a). The average distance to the 5 nearest neighbors is 18km across the 160 core meteorological stations, and 12km when all stations are included. The stations follow the World Meteorological Organization (WMO) standards, use consistent sensors, and are regularly maintained. The measurement data undergo automatic and manual quality controls. From the range of weather parameters available, we select air temperature, relative humidity, atmospheric pressure, sunshine duration, wind speed and direction, wind gusts, and the amount of precipitation, as summarized in Table 1; the table also reports the unique identifier (Short name) consistent with the SwissMetNet naming convention and OGD program (SwissMetNetDocumentation, ). Statistics about parameters measured by different stations are later reported in Table 2 and discussed at the end of the section, highlighting variability across stations.
Figure 1: Visualizations of the placement of PeakWeather stations. Panel 1(a)) distribution of the stations across the Swiss territory; circles denote meteorological stations, while triangles denote rain gauges. Panel 1(b)) a graph obtained using the geographical distance to compute a similarity heuristic for the meteorological stations. Panel 1(c)) vertical coverage of the stations.
@@
temperature tre200s0 Air temperature 2m above the ground. Current value. Celsius
humidity ure200s0 Relative air humidity 2m above the ground. Current value. %
precipitation rre150z0 Precipitation. 10min total. mm
sunshine sre000z0 Sunshine duration. 10min total. min
pressure prestas0 Atmospheric pressure at barometric altitude (QFE). Current value. hPa
wind_speed fkl010z0 Wind speed. 10min mean m/s
wind_gust fkl010z1 Wind gust peak of 1 second 10min maximum. m/s
wind_direction dkl010z0 Wind from direction, measured clockwise from north. 10min mean. degree
Topographic descriptors
Features of the terrain play a crucial role in weather and climate processes across multiple scales and by integrating detailed terrain information, prediction accuracy of data-driven models can be enhanced as demonstrated in Section 5. To this end, PeakWeather includes a high-resolution digital height model (DHM) with a spatial resolution of $50$ meters, capturing the bare ground elevation without vegetation and buildings. The DHM is derived from the DHM25 (swisstopoDHM25, ) product by the Swiss Federal Office of Topography swisstopo; a visualization is provided in Figure 1. In addition, we provide various derived topographical features that represent physical characteristics of the terrain and its surroundings. These features [^1] The implementation of these features has been open-sourced by MeteoSwiss https://github.com/MeteoSwiss/topo-descriptors. are computed considering neighborhoods of $2$ and $10$ kilometers to cover phenomena at various scales and include the topographic position index (TPI) which describes whether a point is above or below the average elevation of its surroundings, the standard deviation (STD) of the elevation, the terrain aspect, the terrain slope, the west-east gradient of the terrain and the south-north gradient of the terrain.
Numerical weather predictions
We include forecasts from ICON-CH1-EPS, one of the NWP models operational at MeteoSwiss since May 2024, as a strong physics-based baseline. ICON-CH1-EPS is a regional, high-resolution, and ensemble setup of the ICOsahedral Non-hydrostatic (ICON zangl2015icon, ) model for Switzerland. It runs every 3 hours and produces an 11-member ensemble: one control run and ten perturbed members. Operating at $\sim$1km grid resolution, it generates forecasts up to 33 hours and the analysis step (lead time 0). PeakWeather includes hourly forecasts of the same variables and extracted at each station’s nearest model grid cell starting from May 14th, 2024.
Table 2: Number of stations per variable and station type, shown as “# in Jan 2017 / # in Mar 2025.” Missing data percentages are computed only for stations that recorded the variable.
@@
Rain gauges 41/38 0/0 124/140 0/0 0/0 0/0 0/0 0/0
Meteo stations 148/149 148/149 138/139 103/131 137/139 149/151 150/151 150/151
Missing values 1.88% 0.94% 3.67% 13.72% 1.39% 1.46% 1.06% 1.08%
@@
One potential source of inconsistencies is sensor or station relocation, which—even in a carefully managed network like SwissMetNet—is sometimes necessary. To ensure transparency, PeakWeather also includes detailed metadata on all sensor relocations which, however, are generally minor: 95% of displacements are under 139 meters, with an average of 104 meters. Moreover, these relocations affect only a minority of stations, up to 17 for relative sunshine duration, and fewer than 10 for most other parameters. Although the stations and data are quality controlled, it should be noted that artifacts may still be present (e.g., due to frozen anemometers), and that the temperature sensors at rain gauges can be of inferior quality, as they are intended primarily to discriminate between solid and liquid precipitation.
@@
Missing data in PeakWeather stems from both temporary measurement gaps and the fact that not all 302 SwissMetNet stations are equipped with the same set of sensors. Sensor configurations also evolve due to operational requirements. To give an accurate overview, Table 2 presents the number of stations recording each variable at the start and end of the dataset (January 2017 and March 2025), along with missing value statistics computed only for stations that have measured each variable at some point. Over the entire period, the overall proportion of missing data is approximately 3%. Missing values are represented as NaNs and can be handled using built-in utilities (e.g., last value carried forward or zero-fill). A binary availability mask is also provided by PeakWeather, allowing users to easily identify and handle invalid observations at each time step, station, and variable level.
Further details and visualizations are provided in Appendices A, B, and C.
@@ 4 Machine learning tasks on PeakWeather
The PeakWeather dataset enables a wide range of machine learning tasks. In this section, we outline several representative examples that can be addressed using the data. While not exhaustive, these tasks demonstrate the dataset’s versatility for both foundational research and practical applications in data-driven weather modeling.
@@
We formalize the dataset as consisting of $N$ multivariate time series, each collected from a different station. Let ${\bm{x}}_{t}^{i}\in\mathbb{R}^{D_{x}}$ denote the $D_{x}$-dimensional observation at time step $t$ from station $i\in{1,\dots,N}$; here, $N$ denotes the total number of weather stations in the dataset and $D_{x}$ the number of possible measured variables (channels) across all stations. Regular and synchronous sampling is assumed. However, as mentioned above, not all stations measure the same weather quantities, and not all sensors are available at all time steps (see Table 2). We account for such data (un)availability by appropriately padding the time series, thus maintaining a tabular representation. Accordingly, an auxiliary binary mask ${\bm{m}}^{i}_{t}\in\{0,1\}^{D_{x}}$ is introduced at each time step $t$ and station $i$ to flag the data availability at the level of the time step, station, and channel. In particular, we set ${\bm{m}}_{t}^{i}[d]=1$ if the $d$-th of the $D_{x}$ channels of the $i$-th station at time step $t$ is valid, and ${\bm{m}}_{t}^{i}[d]=0$ otherwise. For conciseness, we denote by ${\bm{X}}_{t}\in\mathbb{R}^{N\times D_{x}}$ the stack of all observations $\{{\bm{x}}_{t}^{i}\}_{i=1}^{N}$ at time step $t$, by ${\bm{X}}_{t:t+T}\in\mathbb{R}^{T\times N\times D_{x}}$ the observations from time step in $\{t,t+1,\dots,t+T-1\}$ and by ${\bm{X}}_{<t}$ all observations up to time step $t$ (excluded); similarly, ${\bm{M}}_{t}\in\mathbb{R}^{N\times D_{x}}$ denotes the mask across all stations at time step $t$. The PeakWeather library provides functionalities to generate such a mask.
@@
As the time series correlates among themselves, we model the data-generating process as a time-invariant spatiotemporal stochastic process such that
${\bm{x}}_{t}^{i}=p^{i}({\bm{x}}_{t}^{i}\,|\,{\bm{X}}_{<t},{\bm{U}}_{\leq t},{\bm{V}}),$ $\forall i=1,\dots,N,$ (1)
@@
where ${\bm{V}}\in\mathbb{R}^{N\times d_{v}}$ and ${\bm{U}}_{t}\in\mathbb{R}^{N\times d_{u}}$ represent, respectively, static and time-dependent exogenous variables describing additional information related to the measured variables ${\bm{x}}_{t}^{i}$, such as station-specific characteristics, seasonalities, as well as the topographic descriptors from the DHM (see Section 3). Note that the process of Equation 1 generating all time series is not necessarily the same, i.e., $p^{i}\neq p^{j}$ if $i\neq j$, while the assumption of time invariance remains valid.
Relational inductive biases for time series processing
Time series processing can be significantly improved by exploiting the underlying (functional) dependencies among series acquired from spatially distributed stations. Rather than modeling each sequence in isolation, one can incorporate relational information as an inductive bias, allowing models to condition predictions on related time series and reduce the risk of overfitting to local, spurious patterns. These dependencies can be encoded via a graph structure represented by an adjacency matrix ${\bm{A}}\in\{0,1\}^{N\times N}$, where each node corresponds to a station and non-zero entries denote pairwise relationships. When needed, ${\bm{A}}$ may be enriched with edge weights or attributes reflecting the strength or type of interaction (e.g., distance, elevation difference, empirical correlation). This setup is at the core of graph deep learning approaches for time series processing (cini2023graph, ), where message-passing operators like graph neural networks (GNN) are exploited to exchange information across the graph topology. In the context of meteorological data, such relations are often assumed to be time-invariant. However, changes to the sensor network—such as the addition, removal, or relocation of stations—may require considering a dynamic or evolving graph. While spatial proximity naturally induces correlation (i.e., signal smoothness), this assumption breaks down at larger scales or in complex terrains. In alpine regions, for instance, valleys, ridges, and altitude gradients introduce significant variability, making long-range dependencies non-trivial to infer. Particularly in datasets like PeakWeather, this motivates considering relations learned data-driven and tailored to the given task, as discussed in the next experimental section.
@@
for a set of lead times $h=0,\dots,H-1$ and stations $i$ of interest; $W\geq 1$ denotes the time window length used to make the predictions. This can be performed at multiple temporal resolutions and forecast horizons, and can leverage both spatial and temporal dependencies to predict weather conditions across multiple stations simultaneously. This includes using graph-based models to capture correlations between locations that condition predictions on a graph ${\bm{A}}$.

### v2

@@ 1 Introduction
Weather forecasts provide essential information for protecting lives and properties, as well as supporting everyday decision-making. The quality of these forecasts has steadily improved over recent decades, driven by advances in numerical weather prediction (NWP) and high-performance computing (HPC) [1]. NWP involves using mathematical models that simulate the weather based on current conditions of the atmosphere, land surfaces, and oceanic conditions. Both deterministic forecasts and ensembles of simulations are produced, with ensemble forecasts offering a probabilistic view of future states and serving as a critical tool for risk-informed decision-making. However, the complexity of the models, combined with the demand for high-resolution predictions in space and time, makes state-of-the-art NWP computationally expensive, often requiring several hours of runtime on modern HPC systems.
@@
In parallel, the rise of deep learning (DL) has brought remarkable progress to a range of fields, including meteorology [2, 3, 4, 5]. Beyond breakthroughs in model architectures, this progress has been enabled by the availability of long-term, high-quality reanalysis datasets. Reanalyses combine modern NWP models with historical observational data, from ground stations, satellites, radar, ships, and aircraft, to produce a consistent and comprehensive estimate of past atmospheric states. DL models can learn to emulate the underlying physics captured in these massive datasets, enabling accurate and scalable data-driven weather forecasting. One of the key advantages of such models is their efficiency: once trained, they are significantly cheaper to run operationally than classical physics-based NWP. Nonetheless, these models still rely on NWP-generated analyses at inference time to initialize forecasts.
@@
To overcome the need for NWP-based analysis and reanalysis data altogether, Direct Observation Prediction (DOP) methods have been proposed [6]. DOP approaches aim to learn directly from raw observational data, eliminating the need for NWP data during both training and inference. When successfully implemented, DOP models can offer faster, more cost-effective forecasting pipelines and avoid potential biases introduced by physics-based simulations. DOP has shown promising results in short-range weather forecasting, particularly for lead times ranging from minutes to a few hours [7, 8], a domain known as nowcasting. A typical nowcasting application is the extrapolation of radar image sequences to predict imminent precipitation. While DOP has proven effective in short time horizons, NWP systems generally maintain an edge in predictive accuracy beyond a few hours.
Figure 1: PeakWeather station locations across the Swiss territory.
@@
We introduce PeakWeather, a benchmarking dataset designed to advance deep learning research on spatiotemporal modeling. The dataset consists of high-resolution, ground-based meteorological measurements collected from SwissMetNet [9], the network of automatic weather stations operated by the Swiss Federal Office of Meteorology and Climatology (MeteoSwiss). It includes eight key weather variables related to temperature, humidity, precipitation, pressure, sunshine, and wind, recorded every 10 minutes over a period of more than eight years. These observations come from 302 meticulously maintained and quality-validated stations [10], distributed throughout Switzerland, a country characterized by complex topography, shaped by the Swiss Alps, the Central Plateau, and the Jura Mountains; see also Figure 1. In addition to meteorological variables, PeakWeather includes topographic descriptors derived from digital elevation models, capturing the unique characteristics of each station’s surrounding area.
@@
To the best of our knowledge, this is the first openly available dataset that combines long-term, high-frequency ground station data with topographic context and state-of-the-art NWP baselines in complex terrain. The complexity of Switzerland’s topography, combined with seasonalities and time variance of the meteorological quantities, poses unique challenges for modeling spatial and temporal dependencies. These factors make PeakWeather an ideal testbed for advancing data-driven methods in a wide array of machine learning tasks, including—but not limited to—time series forecasting, virtual sensing, graph structure learning, and missing data imputation. To support benchmarking, the dataset also includes ensemble forecasts from a state-of-the-art, high-resolution NWP model that has been operational at MeteoSwiss since May 2024. These forecasts serve as a strong physics-based reference point, enabling direct comparisons with machine learning approaches and promoting research in hybrid and alternative forecasting paradigms.
@@
  • Release of PeakWeather dataset [11]. The dataset contains: (i) A collection of high-quality validated multivariate time series from 302 weather stations comprising eight meteorological variables with a temporal resolution of 10 minutes acquired from January 1st, 2017 to October 13th, 2025; (ii) Topographical descriptors derived from a digital elevation model describing the Swiss terrain over a 50-meter grid; (iii) Numerical weather predictions from a state-of-the-art high-resolution NWP model.
  • A framework-agnostic, easy-to-use and well-documented Python library [12] for seamless interaction with the dataset. The library enables users to easily load, align, and preprocess the data, providing them in ready-to-use tensor formats suitable for deep learning workflows.
  • Evaluation of predictive models for wind forecasting—a task of high practical importance in domains such as renewable energy, aviation, and severe weather management. A range of deep learning architectures, including recurrent neural networks, spatiotemporal graph neural networks, and foundation models, are contrasted against forecasts from a state-of-the-art NWP model. Experiments extend to inductive temperature forecasting and study relevant modeling aspects, such as the use of graph-based modeling and station-specific components. Reported results establish valuable baselines for future research and demonstrate the promising potential of deep learning models as viable alternatives to NWP models in specific weather prediction tasks.
The PeakWeather dataset is released under the CC-BY-4.0 license, with the associated library distributed under the BSD-3-Clause. With PeakWeather, we pursue a twofold objective: first, to facilitate access to high-quality meteorological data from MeteoSwiss, which is being progressively released as Open Government Data starting in 2025 [OGD, 13]; and second, to enable reproducible experimentation that advances foundational machine learning research, meteorology, and sensor-based applications.
@@ 2 Related work
One major category of meteorological datasets consists of gridded reanalysis data produced by NWP models, which assimilate past observations to estimate the weather state. One of the main driving forces behind the recent success of data-driven models is ECMWF Reanalysis v5 [ERA5, 14]. This dataset contains a detailed estimate of the global state of the atmosphere, land surface and oceans since the 1950s, based on the reanalysis of ECMWF’s Integrated Forecasting System (IFS). Derived from ERA5, WeatherBench [15] and its recent successor WeatherBench2 [16] offer standardized benchmarking datasets and tools, to facilitate the evaluation and comparison of machine learning models for weather forecasting. For the task of statistical postprocessing [17], where the goal is to correct systematic errors in NWP forecasts, EUPPBench [18] provides a dataset with aligned IFS forecasts and direct observations for a variety of meteorological variables. So far, however, only coarse-grained NWP data is available in EUPPBench and temporal resolution is limited to 6 hours.
Unlike the datasets above, PeakWeather supports DOP, enabling the development of models that rely solely on past observations and static features, making the final model completely independent from NWP models. While datasets such as WeatherReal [19], Weather2k [20], Monash Weather [21], and StationBench [22] share this DOP focus, PeakWeather distinguishes itself through its unique combination of characteristics not available in other datasets: quality-controlled data with a high spatial density and 10-minute temporal resolution, spanning over eight years across Switzerland’s challenging topography, and physics-based forecasts from an operational NWP model for comparison.
@@ 3 PeakWeather dataset
We introduce PeakWeather, a publicly available dataset hosted on Hugging Face [11]. The dataset is organized into three core components: (i) high-resolution historical surface observations, (ii) static topographic features derived from elevation data, and (iii) ensemble forecasts from an operational NWP model. To facilitate easy access and streamlined preprocessing, we also release an open-source Python library [12] tailored to work seamlessly with the dataset.
@@
Meteorological observations come from the SwissMetNet network, Switzerland’s reference ground-based measurement system for weather and climate monitoring operated by MeteoSwiss. The network in PeakWeather comprises 302 stations in total: 160 standard meteorological stations measuring various weather and climate parameters, complemented by 142 additional stations measuring precipitation. Stations are labeled as either meteorological stations or rain gauges, as shown in Figure 1. The average distance to the 5 nearest neighbors is 18km across the 160 core meteorological stations, and 12km when all stations are included. The stations follow the World Meteorological Organization (WMO) standards, use consistent sensors, and are regularly maintained [10]. The measurement data undergo automatic and manual quality controls. From the range of weather parameters available, we select air temperature, relative humidity, atmospheric pressure, sunshine duration, wind speed and direction, wind gusts, and the amount of precipitation, as summarized in Table 1; the table also reports the unique identifier (Short name) consistent with the SwissMetNet naming convention and Open Government Data (OGD) program [13]. Statistics about parameters measured by different stations are reported in Table 2, highlighting variability across stations.
Topographic descriptors
Features of the terrain play a crucial role in weather and climate processes across multiple scales. By integrating detailed terrain information, the prediction accuracy of data-driven models can be enhanced as demonstrated in Section 5. To this end, PeakWeather includes a high-resolution digital height model (DHM) with a spatial resolution of $50$ meters, capturing the bare ground elevation without vegetation and buildings. The DHM is derived from the DHM25 [23] product by the Swiss Federal Office of Topography swisstopo; a visualization is provided in Figure 1. In addition, we provide various derived topographical features that represent physical characteristics of the terrain and its surroundings. These features [^1] The implementation of these features has been open-sourced by MeteoSwiss https://github.com/MeteoSwiss/topo-descriptors are computed considering neighborhoods of $2$ and $10$ kilometers to cover phenomena at various scales and include the topographic position index (TPI) which describes whether a point is above or below the average elevation of its surroundings, the standard deviation (STD) of the elevation, the terrain aspect, the terrain slope, the west-east and south-north gradients of the terrain; see also Appendix A.
Numerical weather predictions
We include forecasts from ICON-CH1-EPS, one of the NWP models operational at MeteoSwiss since May 2024, as a strong physics-based baseline. ICON-CH1-EPS is a regional, high-resolution, and ensemble setup of the ICOsahedral Non-hydrostatic [ICON, 24] model for Switzerland. It runs every 3 hours and produces an 11-member ensemble: one control run and ten perturbed members. Operating at $\sim$1km grid resolution, it generates forecasts up to 33 hours and the analysis step (lead time 0). PeakWeather includes hourly forecasts of the same variables and extracted at each station’s nearest model grid cell starting from May 14th, 2024. Refer to Appendix B for additional details.
@@
temperature tre200s0 Air temperature 2m above the ground Current value Celsius
humidity ure200s0 Relative air humidity 2m above the ground Current value %
precipitation rre150z0 Precipitation 10min total $m$
sunshine sre000z0 Sunshine duration 10min total min
pressure prestas0 Atmospheric pressure at barometric altitude (QFE) Current value hPa
wind_speed fkl010z0 Wind speed 10min mean m/s
wind_gust fkl010z1 Wind gust peak of 1 second 10min max m/s
wind_direction dkl010z0 Wind from direction, measured clockwise from north. 10min circ. mean degree
Table 2: Number of stations per variable and station type, shown as “# in Jan 2017 / # in Oct 2025.” Missing data percentages are computed only for stations that recorded the variable.
@@
Rain gauges 41/38 0/0 124/141 0/0 0/0 0/0 0/0 0/0
Meteo stations 148/148 148/148 138/138 103/129 137/138 149/150 150/150 150/150
Missing values 1.94% 0.96% 3.51% 13.02% 1.39% 1.46% 1.08% 1.10%
@@
One potential source of inconsistencies is sensor or station relocation, which—even in a carefully managed network like SwissMetNet—is sometimes necessary. To ensure transparency, PeakWeather also includes detailed metadata on all sensor relocations which, however, are generally minor: 95% of displacements are under 110 meters, with an average of 97 meters. Although the stations and data are quality controlled, it should be noted that artifacts may still be present (e.g., due to frozen anemometers), and that the temperature sensors at rain gauges can be of inferior quality, as they are intended to discriminate between solid and liquid precipitation.
@@
Missing data in PeakWeather stems from both temporary measurement gaps and the fact that not all 302 SwissMetNet stations are equipped with the same set of sensors. Sensor configurations also evolve due to operational requirements. To give an accurate overview, Table 2 presents the number of stations recording each variable at the start and end of the dataset (January 2017 and October 2025), along with missing value statistics computed only for stations that have measured each variable at some point. Over the entire period, the overall proportion of missing data is approximately 3%. Figure 6 in the appendix shows the sensor availability over time. Missing values are represented as NaNs and can be handled using built-in utilities (e.g., forward fill or zero-fill). A binary availability mask is also provided by PeakWeather, allowing users to easily identify and handle invalid observations at each time step, station, and variable level.
@@ 4 Machine learning tasks on PeakWeather
The PeakWeather dataset supports a wide range of machine learning tasks. While forecasting is likely the most common application, PeakWeather also enables the investigation of additional tasks and modeling aspects arising from the dataset’s distinctive characteristics. In this section, we outline a set of representative examples that, while not exhaustive, illustrate the dataset’s versatility for both foundational research and practical applications in data-driven weather modeling.
Notation
@@
We formalize the dataset as consisting of $N$ multivariate time series, each collected from a different station. Let ${\bm{x}}_{t}^{i}\in\mathbb{R}^{D_{x}}$ denote the $D_{x}$-dimensional observation at time step $t$ from station $i\in\{1,\dots,N\}$; here, $N$ denotes the total number of weather stations in the dataset and $D_{x}$ the number of possible measured variables (channels) across all stations. Regular and synchronous sampling is assumed. However, as mentioned above, not all stations measure the same weather quantities, and not all sensors are available at all time steps (see Table 2). We account for such data (un)availability by appropriately padding the time series, thereby maintaining a tabular representation. Accordingly, an auxiliary binary mask ${\bm{m}}^{i}_{t}\in\{0,1\}^{D_{x}}$ is introduced at each time step $t$ and station $i$ to flag the data availability at the level of the stations, channels, and time steps. In particular, we set ${\bm{m}}_{t}^{i}[d]=1$ if the $d$-th of the $D_{x}$ channels of the $i$-th station at time step $t$ is valid, and ${\bm{m}}_{t}^{i}[d]=0$ otherwise. For brevity, denote by ${\bm{X}}_{t}\in\mathbb{R}^{N\times D_{x}}$ the stack of all observations $\{{\bm{x}}_{t}^{i}\}_{i=1}^{N}$ at time step $t$, by ${\bm{X}}_{t:t+T}\in\mathbb{R}^{T\times N\times D_{x}}$ the observations from time steps in $\{t,t+1,\dots t+T-1\}$ and by ${\bm{X}}_{<t}$ all observations up to time step $t$; similarly, ${\bm{M}}_{t}\in\mathbb{R}^{N\times D_{x}}$ denotes the mask across all stations at time step $t$. The PeakWeather library provides functionalities to generate such a mask.
Data-generating process
@@
As the time series are correlated with each other, we model the data-generating process as a time-invariant spatiotemporal stochastic process such that
${\bm{x}}_{t}^{i}\sim p^{i}({\bm{x}}_{t}^{i}|{\bm{X}}_{<t},{\bm{U}}_{\leq t},{\bm{V}}),\qquad\forall i=1,\dots,N,$ (1)
@@
where ${\bm{V}}\in\mathbb{R}^{N\times d_{v}}$ and ${\bm{U}}_{t}\in\mathbb{R}^{N\times d_{u}}$ are exogenous variables describing additional information related to the measured variables ${\bm{x}}_{t}^{i}$, such as station-specific characteristics, seasonalities, as well as the topographic descriptors from the DHM (see Section 3). Note that the process of Equation 1 generating all time series is not necessarily the same, i.e., $p^{i}\neq p^{j}$ if $i\neq j$, while the assumption of time invariance remains valid.
@@
for a set of lead times corresponding to $h=0,\dots,H-1$ and stations $i$ of interest; $W\geq 1$ denotes the time window length used to make the predictions. This task can be addressed at multiple temporal resolutions and forecast horizons, and can leverage both spatial and temporal dependencies to predict weather conditions across multiple stations simultaneously.